# Multilingual Pretraining for Pixel Language Models

### v1

@@ 1 Introduction
[…]
@@
In this section, we describe our methodology in detail, including the selection of pretraining languages, the pretraining data creation procedure (§2.1), the model architecture and the pretraining procedure (§2.2).
@@ 2.1 Pretraining Data
Following our motivation to explore multilingual pretraining through a diverse selection of scripts rather than a large range of languages, pixel-m 4 is pretrained on text written in Latin (English), Cyrillic (Ukrainian), Simplified Chinese characters (Chinese), and Devanagari (Hindi). For each script, a corresponding subset of the mC4 (Xue et al., 2021) corpus is rendered into images, following the strategy of rendering two characters per image patch from Lotz et al. (2023). With a sequence length of 529 image patches and a batch size of 256, the model observes approximately 135 billion image patches over 1 million pretraining steps – this is the same total amount of data as the original pixel and pixel-bigrams models. However, pixel-m 4 is trained on an order-or-magnitude more unique samples than pixel-bigrams. This difference is due to the fact that pixel-bigrams was trained by iterating 10 times over the English-only Wikipedia + BookCorpus datasets (Zhu et al., 2015), whereas pixel-m 4 processes each sample in our subset of mC4 only once across the four pretraining languages.
@@ 3 Experimental Setup
This section contains the details of our experiments: [^5] See Appendix for implementation details. §3.1 contains information regarding tasks and benchmarks and §3.2 describes the baselines.
@@
arz uig urd bod ben hin tam tel kir rus ukr
bert-mono 29.1 43.9 31.1 40.7 38.4 87.2 48.6 29.5 73.5 83.8 86.5
pixel-bigrams 38.3 48.6 36.5 36.9 31.7 32.6 39.7 39.9 47.1 37.7 44.4
pixel-m 4 37.5 53.7 41.6 46.3 46.2 78.6 64.5 46.6 62.9 74.7 80.5
@@
Table 1: Text classification results on a selected language subset of the SIB-200 benchmark using macro F1-score. bert-mono indicates that the monolingual bert model varies by language (see §3.2 for details). Best performances are bolded. pixel-m 4 significantly outperforms its English-only-pretrained equivalent pixel-bigrams in almost all non-Latin languages, and pixel-m 4 performs better than monolingual bert models on novel writing systems.
@@ 3.1 Tasks & Benchmarks
We evaluate on the token-level syntactic parsing task of dependency parsing using the Universal Dependencies (UD) benchmark (Nivre et al., 2020; Zeman et al., 2022). We also compare the models using the same three transfer learning settings again: (i) same-script languages seen during pretraining: Latin (English eng, Vietnamese vie), Devanagari (Hindi hin), Han (Chinese zho), and Cyrillic (Ukrainian ukr, Russian rus, Bulgarian bul); (ii) languages in scripts related to at least one pretraining script: Coptic (cop), Japanese (jpn) and Brahmic script languages (Tamil tam, Telugu tel); (iii) languages in scripts unrelated to the pretraining scripts: Arabic abjad (Arabic ara, Urdu urd) and Korean (kor). We report Labeled Attachment Score (LAS) as the evaluation metric.
@@ 3.2 Baselines
We additionally compare pixel-m 4 against four monolingual bert variants: The original English bert (Devlin et al., 2019) primarily for the Latin languages, Chinese bert (Devlin et al., 2019) for Han and Japanese scripts, a Hindi bert (Samuel et al., 2023) for the Brahmic script languages, and a Ukrainian bert (Samuel et al., 2023) for the Cyrillic languages. English bert is also used as a fallback option to evaluate languages that do not belong to any of the pretraining scripts, such as Arabic or Hangul. This allows us to test whether multilingually-pretrained pixel models can match or exceed the cross-lingual transfer capabilities of the tokenizer-based models, not only for Latin scripts but also for others.
@@
Table 2: Dependency parsing results for the selected set of languages in the udp benchmark with LAS. bert-mono indicates that the monolingual bert model varies by language. pixel-m 4 outperforms pixel-bigrams in non-Latin script languages, and it again achieves a better performance than bert-mono on novel scripts.
@@
pixel-m 4 67.3 82.1 80.9 78.5 68.0 79.6 81.6 74.9 75.9
Table 3: NER results by macro-averaged F1-scores. bert-mono is the monolingual bert model varies by language based on their scripts. Overall, pixel-m 4 performs better than pixel-bigrams and bert-mono with an average score of $75.9$.
@@ 4 Results and Discussion
We discuss the results of the downstream task experiments in this section.
@@
Table 2 presents the results on the UDP benchmark. In the pretraining languages, pixel-m 4 significantly improves upon pixel (hin: +3.0, ukr: +9.9, zho: +6.0) except in English (eng: -2.0), which both models have seen in their pretraining. pixel-m 4 outperforms pixel on the languages written in Cyrillic (bul: +2.5, rus: +3.9), which demonstrates improved cross-lingual transfer learning within the same-script languages once again. For the unseen Brahmic languages, pixel-m 4 achieves a slight gain in Telugu (tel: +0.7) and a much larger performance boost in Tamil (tam: +10.7). For the orthographically distant Korean language, pixel-m 4 outperforms pixel-bigrams (kor: +4.2). For the Arabic-script languages, we observe mixed results: In Arabic, the performance drops (ara: -3.5), while we observe a modest gain in Urdu (urd: +0.6). Altogether, multilingually-pretrained pixel-m 4 improves on pixel-bigrams on the dependency parsing task for the unseen languages considering various cross-lingual transfer learning settings. Lastly, our findings on this task is similar to the SIB-200 findings for comparing pixel-m 4 against monolingual bert models: (i) pixel-m 4 achieves a better overall performance than bert-mono in cross-lingual transfer involving writing systems unknown to both; (ii) bert-mono performs better than pixel-m 4 for the pretraining scripts and cross-lingual transfer within the same-script.
Figure 2: Data-efficient learning experiments on the Naamapadam NER benchmark showing the mean test F1 score as a function of training set size in log scale for four Brahmic languages. In each experiment, pixel-m 4 consistently outperforms pixel-bigrams, with the largest relative gains under the smallest data regimes.
@@
Table 3 reports macro-averaged F1 for NER across eight languages. As expected, multilingual pixel pretraining (pixel-m 4 ) outperforms the English-only pixel-bigrams model on every language, raising the average F1 from 73.9 to 75.9. The largest boost is seen in Chinese (zho: +13.5), reflecting that exposure to Chinese during pixel-m4’s pretraining. Other pretraining languages also benefit from multilingual pretraining (eng: +3.9, hin: +1.9). Differently from the other tasks, both pixel-m 4 and pixel-bigrams perform on par in the Brahmic scripts (hin: +1.9, ben: +0.5, tam: +0.1, tel: 0.0): This might be due the larger training sets available in the Naamapadam benchmark. Later, in §5, we show that pixel-m 4 outperforms pixel-bigrams with large margins in low-resource settings. Lastly, +1.2 gain in Korean suggests that pixel-m 4 can transfer visual substructure from unrelated scripts for better entity processing.
The monolingual bert models achieve a better performance than pixel-m 4 for the languages with writing systems known by both models, underscoring that world-knowledge and semantic co-occurrence patterns encoded into specific token entities remain crucial for this semantic task. This is especially the case for English, as both bert and pixel-bigrams are pretrained using exactly the same data. Nonetheless, our findings for the languages in unseen scripts is inline with previous experiments where pixel-m 4 performs better than bert-mono: (ben: +3.1, tam: +0.7, tel: 1.6). These improvements highlight how pixel models can process languages in related scripts directly, avoiding the tokenization failure modes of subword-based models.
@@ 5 Analysis
We perform three different analyses to examine the outcomes of multilingual pretraining, where each subsection covers a different analysis.
@@ 5.2 Word-Level Probing
Figure 4: t-SNE visualization of the outputs for the specified layers. Each row contains visualizations for a particular model, and each column focuses on a particular layer. Each ‘$\bm{\times}$’ marker appear at the centroid of a different pretraining language seen by pixel-m 4 . Both models cluster languages based on their scripts, yet pixel-m 4 clusters some pretraining languages in the later layers.
@@ 5.3 Analyzing Hidden Representations
Similar to Salesky et al. (2023), we visualize the hidden representations learned by both pixel and pixel-m 4 using t-SNE (Van der Maaten and Hinton, 2008). To perform this analysis, we use a subset of SIB-200 (Adelani et al., 2024) including the training splits of 26 languages. We perform t-SNE visualization throughout the model, starting from the convolved input representations (Layer 0) to the output of the last transformer layer (Layer 12). Figure 4 shows t-SNE plots: rows correspond to models, columns to layers, and ‘×’ marks the pixel-m 4 pretraining-language centroids. We observe the same phenomenon for the convolved features as demonstrated in Salesky et al. (2023): Languages which use the same or a related writing script are grouped together. This can be observed for both models, where we can see large clusters for Arabic, Cyrillic and Latin, and Chinese-Japanese language clusters appear next to each other. As we move through in the model layers, we start to see some languages form their own separate clusters by moving away from their script clusters (e.g. Layer 4 and 8). More importantly, in the later layers of pixel-m 4 , we observe that the pretraining languages move away from the rest of the languages that share the same script, and they start to cluster together. This observation demonstrates that pixel-m 4 shifts its focus from visual processing more to the semantics in the later layers. This raises the question of whether pixel-m 4 has learned a semantic representation space shared between different pretraining languages.
@@ 7 Conclusion
pixel-m 4 inherits many of the limitations of its predecessors. First, rendering text using the bigrams strategy leads to increased sequence lengths when a bigram does not fit into single patch. Like Rust et al. (2023) and Lotz et al. (2023), pixel-m 4 cannot generate text. The improvements over pixel-bigrams are also limited for Latin-script languages and also for high-resource settings. Finally, due to our limited compute budget, we pretrained a single pixel-m 4 model on only four languages-each in a different script. Consequently, we have not explored larger or different combinations of languages and scripts, such as additional Latin-script languages (e.g. French, Estonian, Turkish) or right-to-left scripts (e.g. Hebrew, Arabic). We leave these comparisons to future work.
@@ A.2 Implementation Details
[…]
@@
[…]
Weight decay 00
@@
Early stopping patience $20$ $5$ $5$
@@ A.3 LINSPECTOR Results
Figure 18: Word-level probing analysis on linspector for the Polarity task. Each subplot shows a different language; in each, the y-axis represents model accuracies and the x-axis represents layer number of the hidden representations.
Figure 19: Word-level probing analysis on linspector for the Possession task. Each subplot shows a different language; in each, the y-axis represents model accuracies and the x-axis represents layer number of the hidden representations.

### v2

@@ 1 Introduction
  • We present the first multilingually-pretrained general-purpose pixel language model, [^2] Code and models: [Uncaptioned image]ilkerkesen/pixel-m4 trained over four visually and linguistically diverse languages.
@@ 2.1 Pretraining Data
Following our motivation to explore multilingual pretraining through a diverse selection of scripts rather than a large range of languages, pixel-m 4 is pretrained on text written in Latin (English), Cyrillic (Ukrainian), Simplified Chinese characters (Chinese), and Devanagari (Hindi). For each script, a corresponding subset of the mC4 (Xue et al., 2021) corpus is rendered into images, following the strategy of rendering two characters per image patch from Lotz et al. (2023). With a sequence length of 529 image patches and a batch size of 256, the model observes approximately 135 billion image patches over 1 million pretraining steps – this is the same total amount of data as the original pixel and pixel-bigrams models. However, pixel-m 4 is trained on an order-of-magnitude more unique samples than pixel-bigrams. This difference is due to the fact that pixel-bigrams was trained by iterating 10 times over the English-only Wikipedia + BookCorpus datasets (Zhu et al., 2015), whereas pixel-m 4 processes each sample in our subset of mC4 only once across the four pretraining languages. processes each sample only once while still adhering to the same number of update steps.
@@ 3 Experimental Setup
arz uig urd ben bod hin tam tel kir rus ukr
bert-mono 29.1 43.9 31.1 38.4 40.7 87.2 48.6 29.5 73.5 83.8 86.5
canine-s 65.1 53.5 55.5 54.5 39.3 53.4 57.5 44.5 65.5 75.9 74.7
pixel-bigrams 38.3 48.6 36.5 31.7 36.9 32.6 39.7 39.9 47.1 37.7 44.4
pixel-m 4 37.5 53.7 41.6 46.2 46.3 78.6 64.5 46.6 62.9 74.7 80.5
@@
Table 1: Text classification results on a selected language subset of the SIB-200 benchmark using macro F1-score. bert-mono indicates that the monolingual bert model varies by language (see §3.2 for details). Best performances are bolded. pixel-m 4 significantly outperforms its English-only-pretrained equivalent pixel-bigrams in almost all non-Latin languages, and pixel-m 4 performs better than monolingual bert models on novel writing systems. Even though canine-s is pretrained on 104 languages, pixel-m 4 outperforms it on 8 of 23 languages.
@@ 3.1 Tasks & Benchmarks
We evaluate on the token‐level syntactic parsing task of dependency parsing using the Universal Dependencies (UD) benchmark (Nivre et al., 2020; Zeman et al., 2022). We also compare the models using the same three transfer learning settings again: (i) same-script languages seen during pretraining: Latin (English eng, Vietnamese vie), Devanagari (Hindi hin), Han (Chinese zho), and Cyrillic (Ukrainian ukr, Russian rus, Bulgarian bul); (ii) languages in scripts related to at least one pretraining script: Coptic (cop), Japanese (jpn) and Brahmic script languages (Tamil tam, Telugu tel); (iii) languages in scripts unrelated to the pretraining scripts: Arabic abjad (Arabic ara, Urdu urd) and Korean (kor). We report Labeled Attachment Score (LAS) as the evaluation metric.
@@ 3.2 Baselines
We also compare pixel-m 4 against four monolingual bert variants: The original English bert (Devlin et al., 2019) primarily for the Latin languages, Chinese bert (Devlin et al., 2019) for Han and Japanese scripts, a Hindi bert (Samuel et al., 2023) for the Brahmic script languages, and a Ukrainian bert (Samuel et al., 2023) for the Cyrillic languages. English bert is also used as a fallback option to evaluate languages that do not belong to any of the pretraining scripts, such as Arabic or Hangul. This allows us to test whether multilingually-pretrained pixel models can match or exceed the cross-lingual transfer capabilities of the tokenizer-based models, not only for Latin scripts but also for others.
Lastly, we include canine-s (Clark et al., 2022) as a character-level multilingual baseline, which is comparable in size to pixel-m 4 . However, canine-s is pretrained with a subword loss objective on 104 languages using the same multilingual Wikipedia corpus as mBERT (Devlin et al., 2019) whereas pixel-m 4 has only seen four languages in pretraining. Consequently, this difference in pretraining data and languages prevents a fully fair comparison of these models.
@@
Table 2: Dependency parsing results for the selected set of languages in the udp benchmark with LAS. bert-mono indicates that the monolingual bert model varies by language. pixel-m 4 outperforms pixel-bigrams in non-Latin script languages, and it again achieves a better performance than bert-mono on novel scripts, while showing substantial gains over canine-s on every language but one.
@@
pixel-m 4 67.3 82.1 80.9 78.5 68.0 79.6 81.6 74.9 76.6
Table 3: NER results by macro-averaged F1-scores. bert-mono is the monolingual bert model, which varies by language depending on script. Overall, pixel-m 4 performs on par with canine-s, and outperforms pixel-bigrams and bert-mono with an average score of $76.6$.
@@ 4 Results and Discussion
We further compare pixel-m 4 against character-level multilingually-pretrained canine-s. pixel-m 4 outperforms canine-s on 3 out of 4 pretraining languages and 4 out of 5 Brahmic languages, as well as slight gains in kor and uig. These gains are notable given that canine-s was pretrained on all of these languages except arz and bod.
@@
Table 2 presents the results on the UDP benchmark. In the pretraining languages, pixel-m 4 significantly improves upon pixel (hin: +3.0, ukr: +9.9, zho: +6.0) except in English (eng: -2.0), which both models have seen in their pretraining. pixel-m 4 outperforms pixel on the languages written in Cyrillic (bul: +2.5, rus: +3.9), which demonstrates improved cross-lingual transfer learning within the same-script languages once again. For the unseen Brahmic languages, pixel-m 4 achieves a slight gain in Telugu (tel: +0.7) and a much larger performance boost in Tamil (tam: +10.7). For the orthographically distant Korean language, pixel-m 4 outperforms pixel-bigrams (kor: +4.2). For the Arabic‐script languages, we observe mixed results: In Arabic, the performance drops (ara: -3.5), while we observe a modest gain in Urdu (urd: +0.6). Altogether, multilingually-pretrained pixel-m 4 improves on pixel-bigrams on the dependency parsing task for the unseen languages considering various cross-lingual transfer learning settings. Our findings on this task is similar to the SIB-200 findings for comparing pixel-m 4 against monolingual bert models: (i) pixel-m 4 achieves a better overall performance than bert-mono in cross-lingual transfer involving writing systems unknown to both; (ii) bert-mono performs better than pixel-m 4 for the pretraining scripts and cross-lingual transfer within the same-script. Lastly, when compared with canine-s, pixel-m 4 achieves consistently better overall performance, despite the fact that canine-s has been trained on all languages except cop.
Figure 2: Data‐efficient learning experiments on the Naamapadam NER benchmark showing the mean test F1 score as a function of training set size in log scale for four Brahmic languages. In each experiment, pixel-m 4 consistently outperforms pixel-bigrams, with the largest relative gains under the smallest data regimes.
@@
Table 3 reports macro-averaged F1 for NER across eight languages. As expected, multilingual pixel pretraining (pixel-m 4 ) outperforms the English‐only pixel-bigrams model on every language, raising the average F1 from 73.9 to 75.9. The largest boost is seen in Chinese (zho: +13.5), reflecting that exposure to Chinese during pixel-m4’s pretraining. Other pretraining languages also benefit from multilingual pretraining (eng: +3.9, hin: +1.9). Differently from the other tasks, both pixel-m 4 and pixel-bigrams perform on par in the Brahmic scripts (hin: +1.9, ben: +0.5, tam: +0.1, tel: 0.0): This might be due the larger training sets available in the Naamapadam benchmark. Later, in §5, we show that pixel-m 4 outperforms pixel-bigrams with large margins in low-resource settings. Lastly, +1.2 gain in Korean suggests that pixel-m 4 can transfer visual substructure from unrelated scripts for better entity processing.
The monolingual bert models achieve a better performance than pixel-m 4 for the languages with writing systems known by both models, underscoring that world‐knowledge and semantic co‐occurrence patterns encoded into specific token entities remain crucial for this semantic task. This is especially the case for English, as both bert and pixel-bigrams are pretrained using exactly the same data. Nonetheless, our findings for the languages in unseen scripts is inline with previous experiments where pixel-m 4 performs better than bert-mono: (ben: +3.1, tam: +0.7, tel: 1.6). These improvements highlight how pixel models can process languages in related scripts directly, avoiding the tokenization failure modes of subword-based models.
Finally, pixel-m 4 performs on par with canine-s overall, although canine-s was exposed to all downstream task languages during pretraining.
@@ 5.3 Analyzing Hidden Representations
Similar to Salesky et al. (2023), we visualize the hidden representations learned by both pixel-bigrams and pixel-m 4 using t-SNE (Van der Maaten and Hinton, 2008). To perform this analysis, we use a subset of SIB-200 (Adelani et al., 2024) including the training splits of 26 languages. We perform t-SNE visualization throughout the model, starting from the convolved input representations (Layer 0) to the output of the last transformer layer (Layer 12). Figure 4 shows t-SNE plots: rows correspond to models, columns to layers, and ‘×’ marks the pixel-m 4 pretraining-language centroids. We observe the same phenomenon for the convolved features as demonstrated in Salesky et al. (2023): Languages which use the same or a related writing script are grouped together. This can be observed for both models, where we can see large clusters for Arabic, Cyrillic and Latin, and Chinese-Japanese language clusters appear next to each other. As we move through in the model layers, we start to see some languages form their own separate clusters by moving away from their script clusters (e.g. Layer 4 and 8). More importantly, in the later layers of pixel-m 4 , we observe that the pretraining languages move away from the rest of the languages that share the same script, and they start to cluster together. This observation demonstrates that pixel-m 4 shifts its focus from visual processing more to the semantics in the later layers. This raises the question of whether pixel-m 4 has learned a semantic representation space shared between different pretraining languages.
@@ 7 Conclusion
pixel-m 4 inherits many of the limitations of its predecessors. First, rendering text using the bigrams strategy leads to increased sequence lengths when a bigram does not fit into single patch. Like Rust et al. (2023) and Lotz et al. (2023), pixel-m 4 cannot generate text. The improvements over pixel-bigrams are also limited for Latin‐script languages and also for high-resource settings. Due to our limited compute budget, we pretrained a single pixel-m 4 model on only four languages-each in a different script. Consequently, we have not explored larger or different combinations of languages and scripts, such as additional Latin-script languages (e.g. French, Estonian, Turkish) or right-to-left scripts (e.g. Hebrew, Arabic). Finally, the comparison between pixel-m 4 and pixel-bigrams is not entirely fair, as pixel-m 4 was exposed to more data. Nonetheless, pretraining another monolingual model on an equivalent amount of data for such a comparison was not feasible within our computational budget. We leave these comparisons to future work.
@@ A.2 Implementation Details
canine-s. We use the same experimental setup and hyperparameters for canine-s with pixel-m 4 and pixel-bigrams.
Monolingual BERT Models. All models were fine-tuned in 16-bit BrainFloat (Abadi et al., 2016) using AdamW (Kingma and Ba, 2015; Loshchilov and Hutter, 2019) with a maximum learning rate of 5e5510-5 that is warmed up over the first 100 steps and subsequently linearly decayed toward 0. Across all tasks, we fine-tune for at maximum 15,000 steps, while evaluating every 500 steps for dependency parsing and NER, whereas topic classification is evaluated every epoch. Early stopping of 5 evaluation cycles (DP and NER) or 20 epochs with a threshold of 0.0 is implemented. For all tasks and languages, when a separate evaluation split is available, we selected the checkpoint performing best on it and evaluated on the test split. If no separate evaluation split was available, we selected and reported the best performance on the evaluation split. Inputs were truncated or padded to a maximum length of 256 tokens for parsing and classification, and 196 tokens for NER. For parsing and NER, a batch size of 64 is used, while topic classification is trained with batch size 32. We followed Rust et al. (2023) and evaluated dependency parsing using a biaffine parsing head (Dozat and Manning, 2017; Glavaš and Vulić, 2021).
@@
Adam $\varepsilon$ 1e8110-8
Weight decay 0
@@
Early stopping patience 2020 55 55
@@ A.3 LINSPECTOR Results
A.4 Use of AI Assistants
Within this work, we used AI assistants only to generate code for producing the plots.
@@ A.4 Use of AI Assistants
Figure 18: Word‑level probing analysis on linspector for the Polarity task. Each subplot shows a different language; in each, the y‑axis represents model accuracies and the x‑axis represents layer number of the hidden representations.
Figure 19: Word‑level probing analysis on linspector for the Possession task. Each subplot shows a different language; in each, the y‑axis represents model accuracies and the x‑axis represents layer number of the hidden representations.